OpenStudy (anonymous):

medal to how ever helps Find the vertex of the graph y equal x to the power of 2 plus 6 x minus 5

OpenStudy (anonymous):

(-3,-5) (3, 5) (3,14) (-3,-14)

OpenStudy (anonymous):

still don't know?

OpenStudy (anonymous):

Another way is to first find the axis of symmetry with this equation\[X=\frac{ -b }{ 2a }\]

OpenStudy (anonymous):

And plugging in that answer to find y

OpenStudy (cwrw238):

x^2 + 6x - 5 = (x + 3)^2 - 9 - 5 = (x + 3)^2 - 14

OpenStudy (cwrw238):